A given mass of nitrogen is 0.12 dm³ at 60⁰C and 1.01×10⁵ Nm–². Find it's pressure at the same temperature if it's volume is changed to 0.24 dm³.

1
Expert's answer
2022-04-20T11:05:04-0400

P1 = 1.01×105 N m–2

V1 = 0.12 dm3

T1 = T2 = 60°C = constant

P2 = ???

V2 = 0.24 dm3


Solution:

Since the temperature and amount of gas remain unchanged, Boyle's law can be used.

Boyle's gas law can be expressed as: P1V1 = P2V2

To find the final pressure, solve the equation for P2:

P2 = P1V1 / V2

P2 = (1.01×105 N m–2 × 0.12 dm3) / (0.24 dm3) = 5.05×104 N m–2

P2 = 5.05×104 N m–2


Answer: 5.05×104 N m–2